OpenStudy (bossimbacon):

What is the slope of the line? y=-3x+8 slope = @cwrw238

OpenStudy (anonymous):

An equation for a line in slope-intercept form looks like this: y=mx+b, where b is the y-intercept and m is the slope.

OpenStudy (bossimbacon):

@Becki i know the formula and stuff i just need to leave soon and i gotta fish my hw can you help me with a couple questions real fast?

OpenStudy (anonymous):

I don't know how else to help you without flat-out giving you the answer, and I won't do that. Sorry.

OpenStudy (bossimbacon):

alrighty then

OpenStudy (anonymous):

What number is in the place of m in your equation?

OpenStudy (anonymous):

m=30.769 is the answer to the question

OpenStudy (anonymous):

And princed is incorrect.

OpenStudy (bossimbacon):

3 would be the m @becki

OpenStudy (anonymous):

Don't forget your sign. There's a negative in front of the 3.

OpenStudy (bossimbacon):

-3*

OpenStudy (anonymous):

Yes. :)

OpenStudy (bossimbacon):

@Becki so its -3? LOL

OpenStudy (anonymous):

sorry i wanted to write 3.0769

OpenStudy (anonymous):

Yes. -3. @princed, I'm not sure where you got your answer.

OpenStudy (anonymous):

-3 is not correct either

OpenStudy (bossimbacon):

What is the y-intercept of the line? y=-2x-8 y-intercept =

OpenStudy (anonymous):

y-intercept is in the place where b is in the slope-intercept equation.

OpenStudy (bossimbacon):

so then it would be 8?

OpenStudy (anonymous):

sign....?

OpenStudy (bossimbacon):

-8*

OpenStudy (anonymous):

yes. :)

OpenStudy (bossimbacon):

thank you only 3 more?

OpenStudy (anonymous):

sure.

OpenStudy (bossimbacon):

What is the slope, and what is the y-intercept of the line? slope = ________ y-intercept =

OpenStudy (bossimbacon):

and the slope is a fraction help pls

OpenStudy (anonymous):

The slope can be a fraction. Slope just tells you how far up and how far over to make the next point on the line. The numerator is the rise (how far UP or DOWN) and the denominator is the run (how far LEFT or RIGHT).

OpenStudy (bossimbacon):

What is the slope, and what is the y-intercept of the line? y=5/4x+9 slope = ________ y-intercept =

OpenStudy (bossimbacon):

so then the slope would be 5/4?

OpenStudy (bossimbacon):

and the y intercept would be 9?

OpenStudy (anonymous):

YES! to both. :)

OpenStudy (bossimbacon):

What is the slope-intercept form of this equation? 3x + 9y = 18 A. y=-1/3x+6 B. y=-1/3x+2 C. y=1/3x+6 D. y=2/3x+2

OpenStudy (anonymous):

Do you know how to solve for a variable? That's all you have to do to put an equation in slope-intercept form. Solve the equation for y.

OpenStudy (bossimbacon):

actually i got it, i got b

OpenStudy (texaschic101):

its B

OpenStudy (anonymous):

Yes.

OpenStudy (bossimbacon):

thank you so much @becki and @texaschic101

OpenStudy (bossimbacon):

saved my life

OpenStudy (anonymous):

no problem. :)